# GRIL-Calib: Targetless Ground Robot IMU-LiDAR Extrinsic Calibration Method using Ground Plane Motion Constraints

### v1

@@ VII CONCLUSIONS
In this paper, we present GRIL-Calib: a novel targetless IMU-LiDAR extrinsic calibration framework that mainly focuses on the ground robot. Our framework aimed to address the challenge of obtaining accurate 6-DoF extrinsic parameters including unobservable direction when the robot’s motion is limited to a planar motion. To achieve this, we proposed the Ground Plane Motion (GPM) constraint to utilize ground observation, which is easily accessible for the ground robots. We also reformulate the calibration as a Single Optimization (SO) problem to take advantage of the correlation of the residuals. Additionally, we proposed ground plane residual to reduce the z-drift of the LiDAR odometry, thereby improving the accuracy of the calibration. Finally, we demonstrated the effectiveness and reliability of our GRIL-Calib by applying it to three public datasets and varying some initial calibration parameters. We acknowledge that our algorithm’s performance might be affected by uneven ground. Future works should focus on calibration for such terrains.

### v2

@@ VII CONCLUSIONS
In this paper, we present GRIL-Calib: a novel targetless IMU-LiDAR extrinsic calibration framework that mainly focuses on the ground robot. Our framework aimed to address the challenge of obtaining accurate 6-DoF extrinsic parameters including unobservable direction when the robot’s motion is limited to a planar motion. To achieve this, we proposed the Ground Plane Motion (GPM) constraint to utilize ground observation, which is easily accessible for the ground robots. We also reformulate the calibration as a Single Optimization (SO) problem to take advantage of the correlation of the residuals. Additionally, we proposed ground plane residual to reduce the z-drift of the LiDAR odometry, thereby improving the accuracy of the calibration. Finally, we demonstrated the effectiveness and reliability of our GRIL-Calib by applying it to various datasets. We acknowledge that our algorithm’s performance might be affected by uneven ground. Future works should focus on calibration for such terrains.